# Which Demographic Features Are Relevant for Individual Fairness Evaluation of U.S. Recidivism Risk Assessment Tools?

### v1

@@ Abstract.
Despite its U.S. constitutional foundation, the technical “individual fairness” criterion has not been operationalized in state or federal statutes/regulations. We conduct a human subjects experiment to address this gap, evaluating which demographic features are relevant for individual fairness evaluation of recidivism risk assessment (RRA) tools. Our analyses conclude that the individual similarity function should consider age and sex, but it should ignore race.
@@ 1. Introduction
However, technical fairness criteria might be hard to enforce without legal incentives. Nguyen et al. (2025) reviewed primary sources of U.S. state/federal law and found that while three major technical fairness criteria (procedural, group, and individual fairness) can be traced back to judicial interpretation of the “Equal Protection” and “Due Process” constitutional clauses, individual fairness has yet been operationalized in specific statutes or regulations. To facilitate the development of more concrete legal criteria on individual fairness, we conduct an original human subjects experiment to determine whether individual fairness should consider or ignore demographic features like race, sex, or age in the individual similarity function (used to identify pairs of highly similar individual, so that the outcomes within each pair of similar defendants can be compared to check if they are similar enough to be “fair”).111 We store Supplementary Materials (e.g., survey form, data, Python code) in a Google Drive folder: https://drive.google.com/drive/folders/1OHyRIQ21ECEZwD2i1xE_7Ahzhqt9w1IO?usp=drive_link
@@ 2. Methodology – Human Study
Once we decide on a simple type of similarity function, we can apply it to a set of features that are surely below the individual fairness scrutiny threshold, e.g., criminal history features, and use this similarity function to run the “badness” score estimation procedure above to get a baseline “badness” score. Then we can evaluate a new similarity function that has the same form but additionally takes into account a feature in dispute, e.g., race, and repeat the procedure to get a “badness” score for this new similarity function. At the final step, if the “badness” score of the similarity function with race is (with statistical significance) higher than the baseline similarity function (without any disputed features), we may conclude that based on humans’ fairness perception of the recidivism risk assessment task, race should be ignored in the individual similarity function, thus above the individual fairness scrutiny threshold.
@@ 3.1. Criminal Justice Dataset
We draw defendant pairs from a real dataset, Client Legal Utility Engine (CLUE), with over 4.4 million District Court and 700,000 Circuit Court cases scraped from the Maryland Judiciary Case Search. [^2] https://casesearch.courts.state.md.us/casesearch/ CLUE contains criminal history and demographic features, e.g. race, sex, and age (at the time of arrest).
@@ 3.2. Main Experiment
Therefore, we conducted a between-subjects experiment on Qualtrics (with participants recruited from Prolific). To enhance data quality and fair compensation, we apply the following participant filters: based in the U.S., English proficiency, Prolific approval rate of at least $98\%$, and Bachelor’s degree. We also provide fairness voting bonuses to incentivize participants to align their fairness evaluation with common sense and optional rationale bonuses to incentivize them to think carefully about each fairness judgment. Furthermore, we filter out responses from participants who fail two or more attention checks. The median completion time among participants who passed our attention checks is $18.5$ minutes. We pay each participant \$$\$3.7$, corresponding to $12$ USD per hour.
@@ 4.2. Quantitative analysis: no significant fairness judgement differences across conditions
Population Fair “Fair” Rating Rate (%)
@@
Table 1. The “Fair” rating rate for each condition
@@
More statistical tests, including OLS Regression, Logistic Regression, and Mixed Effect Linear Regression, also showed no significant difference across our four experimental conditions. Therefore, rigorous quantitative analysis shows no evidence that including a demographic feature (race, sex, or age) would worsen people’s individual fairness judgment (by making them more likely to select “‘Fair”’ even though similar defendants got opposite outcomes). However, this result suggests another potential effect that the multiple-choice questions alone may not capture. Some participants might have internalized certain social norms to not consider certain demographic features (e.g., race) at all when making fairness judgments. We conduct qualitative analysis to investigate whether the quantitative result is because people simply refuse to consider each demographic feature or because each demographic feature, when thoroughly considered, does not worsen individual fairness judgment.
@@
We qualitatively code the optional rationales after each “Fair” vs. “Unfair” vote in the treatment conditions, to check whether the lack of quantitative evidence (that demographic features worsen fairness judgement) is simply because participants ignore the respective demographic feature, or due to more nuanced considerations.
@@
Second, for Age, the biggest cluster is (Fair $\times$ Impact), which is also the most concerning cluster because rationales here show that the demographic feature worsens the quality of the pairwise individual similarity function, by making participants less likely to recognize that the two defendants in the same pair are highly similar and therefore less likely to select the “Unfair” vote. The main recurring theme in this cluster is that it’s fair for a younger person to receive a higher risk prediction, e.g., “The younger person may not be as mature as the older one, but also they have a lot more time in years cause they’re young to commit more crimes.” For the next biggest cluster, a frequent argument in the (Unfair $\times$ Impact) cluster is that being in an older age group means a defendant got enough time to rehabilitate, so their slightly worse criminal record is not enough to justify their higher predicted risk, e.g., “Both should have the same risk. Person 1 does not have a violent conviction. Though person 2 has a prior conviction their age must be considered, this could have been from over 20 years ago.”
It is tempting to conclude that Age worsens the quality of the pairwise individual similarity function more than the other two demographic features because Age is the only feature with a majority (Fair $\times$ Impact) cluster. However, combining this observation and the lack of evidence for cross-condition significant difference, we conclude that Age does not necessarily make more people more likely to choose the “Fair” judgment, but this feature prompts people who choose the “Fair” judgment to think out loud more (by writing a detailed rationale citing age) to justify their choice – a constructive impact.
@@ 5. Discussion: Which Demographic Features Should the Individual Similarity Function Ignore?
First, age belongs to the “rational basis” scrutiny category, which includes the least judicially scrutinized features in demographic classification by the government (U.S. Supreme Court, 1976), so it is more likely than race or sex to be justifiable as part of individual fairness evaluation.
Second, due to the continuous nature of age, a small difference in age, if close to a legal threshold, can yield significant differences in the standard of proof a person is subject to and, therefore, to the conviction outcome. For example, while American adults have a constitutional right to a jury trial, the U.S. Supreme Court ruled in McKeiver v. Pennsylvania (1971) that this constitutional right does not extend to juvenile defendants (U.S. Supreme Court, 1971), which might imply that the standard of proof for an adult v. a juvenile might differ due to the difference in the number of fact finders between a bench trial and a jury trial. A more specific example regarding the age-based difference in the standard of proof is the case In re Mitchell P (1978), where the California Supreme Court ruled that “uncorroborated testimony of an accomplice” is not forbidden as evidence in California juvenile courts (California Supreme Court, 1978), even though California criminal courts (for adults) forbid such evidence, as pointed out by McGoldrick Jr (1978). These subtle differences in standard of proof may make two highly similar-aged defendants (e.g., one old juvenile and one early adult) receive very different criminal history features (such as past convictions). If a pairwise individual similarity function does not take into account a defendant’s age but only the criminal history features, this function might assume that those two highly similar-aged defendants are very different in the input space, and thus make no (un-)fairness alerts on their disparate risk scores.
@@
The main arguments for including sex, typically an “intermediate scrutiny” feature, into the individual similarity function are that first, our empirical study shows that including sex does not significantly worsen people’s fairness judgement, and second, the violence prevalence argument for age also applies to sex (males are often more capable of violence than females). Therefore, the individual similarity function should also include sex.
@@
Decided by the U.S. Supreme Court as a “strict scrutiny” feature (U.S. Supreme Court, 1944), [^5] Korematsu v. United States. U.S. Supreme Court. 323 U.S. 214 (1944): “all legal restrictions which curtail the civil rights of a single racial group are immediately suspect. […] courts must subject them to the most rigid scrutiny.” race is a nuanced case. On the one hand, past studies have shown correlations between race and several features that are related to violence prevalence. For example, Williams et al. (2007) show that among urban juveniles, race has a significant correlation with ‘initiation of [under-aged] alcohol use’ while does not. Between two minority groups of juveniles (Black vs. Hispanic), Reingle et al. (2012) find significantly higher longitudinal level of aggression among Black juveniles. If one believes that defendants with similar prevalence of violence should get similar recidivism prediction, then one might argue that race should be used in the individual similarity function as one of the proxies to estimate similarity in terms of violence prevalence.
@@ Conclusion
We find no statistical evidence that race, sex, nor age worsens laypeople’s individual fairness perception in the RRA context, but qualitative analysis shows that age does impact and make their fairness judgement more nuanced. Our follow-up legal and policy analyses conclude that individual fairness evaluation of RRA tools should consider age and sex, but it should ignore race.

### v2

@@ Abstract.
Despite its constitutional relevance, the technical “individual fairness” criterion has not been operationalized in U.S. state or federal statutes/regulations. We conduct a human subjects experiment to address this gap, evaluating which demographic features are relevant for individual fairness evaluation of recidivism risk assessment (RRA) tools. Our analyses conclude that the individual similarity function should consider age and sex, but it should ignore race.
@@ 1. Introduction
However, technical fairness criteria might be hard to enforce without legal incentives. Nguyen et al. (2025) reviewed primary sources of U.S. state/federal law and found that while three major technical fairness criteria (procedural, group, and individual fairness) can be traced back to judicial interpretation of the U.S. “Equal Protection” and “Due Process” constitutional clauses, individual fairness has yet been operationalized in specific statutes or regulations. To facilitate the development of more concrete legal criteria on individual fairness, we conduct an original human subjects experiment to determine whether individual fairness should consider or ignore demographic features like race, sex, or age in the individual similarity function (to identify pairs of highly similar individual, so that the outcomes within each pair of similar defendants can be compared to check if these outcomes are similar enough to be “fair”).111 We store Supplementary Materials (e.g., survey form, data, Python code) in a Google Drive folder: https://drive.google.com/drive/folders/1OHyRIQ21ECEZwD2i1xE_7Ahzhqt9w1IO?usp=drive_link
@@ 2. Methodology – Human Study
Once we decide on a simple type of similarity function, we can apply it to a set of features that should certainly be included in the similarity function, e.g., criminal history features, and use this similarity function to run the “badness” score estimation procedure above to get a baseline “badness” score. Then we can evaluate a new similarity function that has the same form but additionally takes into account a feature in dispute, e.g., race, and repeat the procedure to get a “badness” score for this new similarity function. At the final step, if the “badness” score of the similarity function with race is (with statistical significance) higher than the baseline similarity function (without any disputed features), we may conclude that based on humans’ fairness perception of the recidivism risk assessment task, race should be ignored, i.e., not included, in the individual similarity function when evaluating individual fairness.
@@ 3.1. Criminal Justice Dataset
We draw defendant pairs from the Client Legal Utility Engine (CLUE) dataset with 4.4 million District Court and 700,000 Circuit Court cases scraped from the Maryland Judiciary Case Search. [^2] https://casesearch.courts.state.md.us/casesearch/ It contains criminal history features (e.g., arrests and convictions) and demographic features, e.g. race, sex, and age (at last arrest).
@@ 3.2. Main Experiment
Therefore, we conducted a between-subjects experiment on Qualtrics (with participants recruited from Prolific). To enhance ecological validity and data quality, we apply the following participant filters: based in the U.S., English proficiency, Prolific approval rate of at least $98\%$, and Bachelor’s degree. We also provide fairness voting bonuses to incentivize participants to align their fairness evaluation with common sense and optional rationale bonuses to incentivize them to think carefully about each fairness judgment. Furthermore, we filter out responses from participants who fail two or more attention checks. The median completion time among participants who passed our attention checks is $18.5$ minutes. We pay each participant \$$\$3.7$, corresponding to $12$ USD per hour.
@@ 4.2. Quantitative analysis: no significant fairness judgement differences across conditions
Ratings Fair “Fair” Rating Rate (%)
@@
Table 1. The “Fair” rating rate per condition. Each of $N=234$ participants who pass all attention checks gives $14$ ratings.
@@
More statistical tests, including OLS Regression, Logistic Regression, and Mixed Effect Linear Regression, also showed no significant difference across our four experimental conditions. Therefore, rigorous quantitative analysis shows no evidence that including a demographic feature (race, sex, or age) would worsen people’s individual fairness judgment (by making them more likely to select “‘Fair”’ even though similar defendants got opposite outcomes). However, this result suggests another potential effect that the multiple-choice questions alone may not capture. Some participants might have internalized certain social norms to not consider certain demographic features (e.g., race) at all when making fairness judgments. We conduct qualitative analysis to investigate whether the quantitative result is because people simply refuse to consider each demographic feature or because each demographic feature, when thoroughly considered, does not worsen people’s individual fairness judgment.
@@
We qualitatively code the optional rationales after each “Fair” vs. “Unfair” vote in the treatment conditions, to check whether the lack of quantitative evidence (that demographic features worsen fairness judgement) is simply because participants ignore each respective demographic feature, or due to more nuanced considerations.
@@
Second, for Age, the biggest cluster is (Fair $\times$ Impact), which is the most seemingly concerning cluster because rationales here show that the demographic feature worsens the quality of the pairwise individual similarity function, by making participants less likely to recognize that the two defendants in the same pair are highly similar and therefore less likely to select the “Unfair” vote. The main recurring theme in this cluster is that it’s fair for a younger person to receive a higher risk prediction, e.g., “The younger person may not be as mature as the older one, but also they have a lot more time in years cause they’re young to commit more crimes.” For the next biggest cluster, a frequent argument in the (Unfair $\times$ Impact) cluster is that being in an older age group means a defendant got enough time to rehabilitate, so their slightly worse criminal record is not enough to justify their higher predicted risk, e.g., “Both should have the same risk. Person 1 does not have a violent conviction. Though person 2 has a prior conviction their age must be considered, this could have been from over 20 years ago.”
It is tempting to conclude that Age worsens the quality of the pairwise individual similarity function more than the other two demographic features because Age is the only feature with a majority (Fair $\times$ Impact) cluster. However, combining this observation and the lack of evidence for cross-condition significant difference, we conclude that Age does not necessarily make more people more likely to choose the “Fair” judgment, but this feature prompts people who choose the “Fair” judgment to think out loud more (by writing a detailed rationale citing age) to justify their choice – a constructive impact to enable more transparent decisions.
@@ 5. Discussion: Which Demographic Features Should the Individual Similarity Function Ignore?
First, age belongs to the “rational basis” category (U.S. Supreme Court, 1976b), [^5] Massachusetts Bd. of Retirement v. Murgia (1976): “We [U.S. Supreme Court] turn then to examine this state [age-based] classification under the rational basis standard.” which includes the least judicially scrutinized features in demographic classification by the government, so it is more likely than race or sex to be justifiable as part of individual fairness evaluation.
Second, due to the continuous nature of age, a small difference in age, if close to a legal threshold, can yield significant differences in the standard of proof a person is subject to and, therefore, to the conviction outcome. For example, while American adults have a constitutional right to a jury trial, the U.S. Supreme Court ruled in McKeiver v. Pennsylvania (1971) that this constitutional right does not extend to juvenile defendants (U.S. Supreme Court, 1971), which might imply that the standard of proof for an adult versus a juvenile might differ due to the difference in the number of fact finders between a bench trial and a jury trial. A more specific example regarding the age-based difference in the standard of proof is the case In re Mitchell P (1978), where the California Supreme Court ruled that “uncorroborated testimony of an accomplice” is not forbidden as evidence in California juvenile courts (California Supreme Court, 1978), even though California criminal courts (for adults) forbid such evidence, as pointed out by McGoldrick Jr (1978). These subtle differences in standard of proof may make two highly similar-aged defendants (e.g., one old juvenile and one early adult) receive very different criminal history features (such as past convictions). If a pairwise individual similarity function does not take into account a defendant’s age but only the criminal history features, this function might assume that those two highly similar-aged defendants are very different in the input space, and thus make no (un-)fairness alerts on their disparate risk scores.
@@
The main arguments for including sex, typically an “intermediate scrutiny” feature (U.S. Supreme Court, 1976a), [^6] Craig v. Boren (1976): “classifications by gender must serve important governmental objectives and must be substantially related to achievement of those objectives.” into the individual similarity function are that first, our empirical study shows that including sex does not significantly worsen people’s fairness judgement, and second, the violence prevalence argument for age also applies to sex (males are often more capable of violence than females). Therefore, the individual similarity function should also include sex.
@@
Decided by the U.S. Supreme Court as a “strict scrutiny” feature (U.S. Supreme Court, 1944), [^7] Korematsu v. United States (1944): “all legal restrictions which curtail the civil rights of a single racial group are immediately suspect. […] courts must subject them to the most rigid scrutiny.” race is a nuanced case. On the one hand, past studies have shown correlations between race and several features that are related to violence prevalence. For example, Williams et al. (2007) show that among urban juveniles, race has a significant correlation with ‘initiation of [under-aged] alcohol use’ while sex does not. Between two minority groups of juveniles (Black vs. Hispanic), Reingle et al. (2012) find significantly higher longitudinal level of aggression among Black juveniles. If one believes that defendants with similar prevalence of violence should get similar recidivism prediction, then one might argue that race should be used in the individual similarity function as one of the proxies to estimate similarity in terms of violence prevalence.
@@ Conclusion
We find no statistical evidence that race, sex, or age worsens laypeople’s individual fairness perception in the RRA context, but qualitative analysis shows that age does impact and make their fairness judgement more nuanced. Our follow-up legal and policy analyses conclude that individual fairness evaluation of RRA tools should consider age and sex, but it should ignore race.